40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Dec 31 19:33:47 2013, tara, Summary, NoiseBudget, beat measurement beat_2013_12_24.pngTTFSS.pngRCAV_OLG.pngACAV_OLGTF.png
    Reply  Wed Jan 15 10:56:08 2014, Evan, Summary, NoiseBudget, Beat measurement: residual short_nb_residual.pdfshort_nb_residual.fig
    Reply  Fri Jan 17 12:46:29 2014, Evan, Summary, NoiseBudget, Noise budget fitting: code fork nb_beat_short.pdf
       Reply  Mon Jan 20 19:20:41 2014, Evan, Summary, NoiseBudget, Noise budget fitting cost_func.pdfnb_beat_short.pdfnb_beat_resid.pdf
          Reply  Wed Jan 22 16:18:37 2014, Evan, Summary, NoiseBudget, Noise budget fitting: need uncertainties 
             Reply  Wed Jan 29 05:44:51 2014, tara, Summary, NoiseBudget, Noise budget fitting: need uncertainties 
                Reply  Wed Jan 29 21:36:53 2014, tara, Summary, NoiseBudget, error in spacer brownian noise  spacerBR_comsol.pngspacerBR_comsol.fig
             Reply  Fri Jan 31 00:44:47 2014, Evan, Summary, NoiseBudget, Noise budget fitting: result 
                Reply  Tue Feb 4 00:31:32 2014, Evan, Summary, NoiseBudget, Noise budget fitting: result 
                Reply  Sun Feb 9 20:37:38 2014, Evan, Summary, NoiseBudget, Noise budget fitting: result lossfit.pdf
Message ID: 1393     Entry time: Tue Dec 31 19:33:47 2013     Reply to this: 1395   1396
Author: tara 
Type: Summary 
Category: NoiseBudget 
Subject: beat measurement 

I got a chance to measure beat measurement. The noise budget is updated and contains all dominant noise traces.

 

== Beat measurement ==

beat_2013_12_24.png

1) at DC to 10Hz, the contribution is mostly from RIN driven Photothermal noise and a bit of seismic noise, a small peaks around 10Hz is probably from the stack, not the cavity sagging. The hump from DC to ~ 50Hz disappear when it is quiet. I think it is mostly scattered light associated with the seismic noise, not displacement noise due to the vibration.

2) 10Hz to 1kHz is pretty much Coating Brownian noise.

3) At 1kHz and above, it is PLL readout noise and residual frequency noise from the laser, where the gain cannot suppress enough noise. This is mostly from ACAV. The residual frequency noise = free running noise / (1+ OLGTF). The measurement of the open loop gain is explained below.

 

==TTFSS Loop characterization==

The OLG TF of TTFSS is measured up to 10MHz and compared with the calculation. The schematic explaning how TTFSS actuates on the laser is shown below.

TTFSS.png

The freqeuncy discriminator can be measured from the slope of the error signal (from Common out1) while scanning the laser. For RCAV Dv = 1/ (194 kHz/V) and 1/(164kHz/V) for ACAV. with 1mW input power.

The adjustable gain stage can be tuned by turning the dial knob. At 400, gain=1, and the gain changes by 10dB with every 250click.

The PZT actuator has a gain of  4.5MHz/V (measured), and the EOM actuator is 15mRad/V (or 15mHz/f  Hz/V) (taken from the spec sheet).

OLG measurement is taken:  RCAV OLG is measured and plotted against the theoretical approximation, see the below figure.

RCAV_OLG.png

above: RCAV OLG TF. Note: The calculation and the measurement do not include the integrator with corner frequency at 4.6kHz.

 

There are some problems with ACAV loop and I could not increase the gain up as much as it used to be and the UGF is around only 200kHz , but the measurement matches the calculation. Right now RCAV servo has a better loop performance.

ACAV_OLGTF.png

 The calculated OLG TF trace(green) should go down at 1MHz or above because of the opamps' bandwidth. I used ideal Op Amps in the simulation because I don't have some op amps in my liso library. I'll see if I can fix it.

ELOG V3.1.3-